OpenStudy (anonymous):

You roll two dice. What is the probability that the sum of the dice is greater than 6 and one die shows a 2? A 6 X 6 table of dice outcomes will help you to answer this question. a. 3/4 b. 1/9 c.1/18 d.4/21

OpenStudy (haseeb96):

O(S)= 36 and the one dice show 2 so other dice will be 5 and 6 i.e.(5,2),(6,2) so O(A)=2 probability = O(A)/O(S) =2 /36 =1/18 is the correct answer

OpenStudy (anonymous):

it was not

OpenStudy (haseeb96):

what ?

OpenStudy (haseeb96):

why ?

OpenStudy (amistre64):

for starters, it was a direct answer ....

OpenStudy (haseeb96):

okay how can i tell him ?

OpenStudy (amistre64):

ask him to make a table

OpenStudy (anonymous):

thats what i did and i got 1/9

OpenStudy (amistre64):

lets review the table then 1 2 3 4 5 6 1 2 3 4 5 6 7 2 3 4 5 6 7 8 3 4 5 6 7 8 9 4 5 6 1+2+3+4+5+6 results is what im getting

OpenStudy (amistre64):

and one shows a 2 .... lets factor that into the setup

OpenStudy (amistre64):

1 2 3 4 5 6 1 2 3 4 5 6 7 2 3 4 5 6 7 8 <-- 2 3 4 5 6 7 8 9 4 5 6 7 8 9 10 5 6 7 8 9 10 11 6 7 8 9 10 11 12 ^ 2

OpenStudy (haseeb96):

lets look at here there is one dice in which only 2 will come and the other dice the number greater than 4 will come because according to yur question when the sum of two dice (5,2) =7 is greater than 6 (6,2)=8 is greater than 8 so only the two events are there so we write it as O(A)=2 and the sample space is 36 ,there are 6 posiblities in every dice so 6^2 or 6 x 6 =36 O(S)=36 probability = O(A) /O(S) now just put the value into the formula and get the answer

OpenStudy (amistre64):

i see 4 out of 36

OpenStudy (amistre64):

2,5 and 2,6 are also greater than 6

OpenStudy (amistre64):

so yes, 4/36 = 1/9

OpenStudy (anonymous):

see now thats how you help some one

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

hasbees concept was sound, but just missing a few of the outcomes :)